OpenStudy (anonymous):

Please help, precalc test tomorrow! Question is find all solutions on interval 0 less than or equal to theta less than or equal to 2 pi for cos(2theta)+1=0 I'm confused on how to get 2theta.....

OpenStudy (danjs):

do you remember the cosine of (2 *angle) = cos(2x) = cos2(x) – sin2(x) = 1 – 2sin2(x) = 2cos2(x) – 1 those "double angle" identities they probably derived for you in trig awhile back

OpenStudy (anonymous):

I've never taken trig

OpenStudy (danjs):

ok, well they are listed in trig identity tables... sorry the past there was bad \[\cos(2x) = \cos^2(x)-\sin^2(x) = 1 - 2*\sin^2(x) = 2*\cos^2(x) - 1\]

OpenStudy (danjs):

algebra 2 class you may have seen it, alg2/trig

OpenStudy (danjs):

you can use one of those, or just see that cos(2*angle) = -1 the cosine value is -1 for a certain angle in the circle, the -x direction , 180 deg, pi

OpenStudy (danjs):

cos(u) = -1 u = pi so 2*angle has to be pi angle is then pi/2

OpenStudy (anonymous):

Ok.... so how did you get from 2cos^2theta-1 = cos(2theta)=-1?